OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. x2 – 10x = 15 pls help ty

OpenStudy (hoblos):

x2 – 10x = 15 x2 -10x +25 - 25 = 15 (x-5)^2 = 40

OpenStudy (anonymous):

ty for the help much appreciated

OpenStudy (hoblos):

my pleasure XD